Fermi National Laboratory


Accelerator Update for June 8, 2005 - June 10, 2005

Wednesday June 8
The day shift began with Operations monitoring the Recycler stash, with the Antiproton Source (Pbar) stacking, and with the MiniBooNE and NuMI experiments taking beam. The Collider Detector at Fermilab (CDF) and D-Zero (D0) experimenters accessed their collision halls.

At 8:50 AM, Operations reported that a Booster kicker (MKS04) had stopped firing. An operator informed Booster experts who bypassed the kicker.

CDF came out of access at 9:40 AM and D0 came out at 10:03 AM.

At 10:48 AM, Operations reported a beam problem in the Main Injector (MI). Operators worked on the Booster and MI coalescing.

Tevatron (TeV) experts began a study period around 11:13 AM.

At 12:46 PM, Operations reported a problem with the Pbar beam cooling systems. Operators stopped stacking while experts investigated.

At 1:02 PM, the decision was made to stop the TeV studies and begin preparations for going into shot setup to use the stack due to the Pbar cooling problems.

Operations put the TeV through a dry squeeze at 1:37 PM.

At 2:08 PM, Operations reported that beginning shot setup was on hold due to problems with a Pbar Debuncher power supply (D:V901). An expert had it fixed within ten minutes.

Operations established combination store 4198 at 4:15 PM with an initial luminosity of 86.47E30.

Operations prepared Pbar for an access at 4:19 PM. The experts came out of access at 5:19 PM.

At 6:13 PM, Pbar experts reported that a transformer contactor for an Accumulator power supply (A:IB) wasn't operating properly. The experts began repairs.

At 6:48 PM, Operations reported that NuMI had tripped off due to beam position problems for the first of four times in ninety minutes. It reset okay each time.

One of the Pbar experts working on A:IB reported at 8:06 PM that they had replaced a 480 volt contactor. Pbar resumed stacking at 8:36 PM.

At 9:22 PM, Operations reported the first of many Booster trips due to beam positions. Experts and operators began tuning the beam.

Thursday June 9
The midnight shift began with Operations monitoring store 4198 and a very small stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

At 2:34 AM, Operations reported that the I- Source started sparking and then tripped off. Operators investigated and changed the voltage command module. Beam resumed at 2:53 AM.

Operations put the Recycler in shot setup to stash antiprotons at 3:10 AM. The transfers ended at 3:53 AM and Pbar resumed stacking.

At 11:22 AM, the Booster chopper tripped off due to sparking. Experts investigated and found a bad cable. Booster beam resumed around 12:30 PM.

At 4:15 PM, the Booster beam permit dropped for the first of many times due to beam positions. Experts lowered the intensity, tuned the machine, and then returned the Booster to normal operations at 5:23 PM.

At 5:52 PM, a Booster kicker (MKS08) tripped off. Operator investigated. Experts got involved and found a PFN problem. Kicker technicians were called in.

Operators retimed the other Booster kickers at 6:30 PM, so that stacking could resume as well as sending beam to NuMI.

At 7:54 PM, the NuMI beam permit dropped due to a MI62 kicker problem. Operators called in an expert.

Operations reported at 8:39 PM that the Booster was back up and running.

Operations terminated store 4198 at 8:59 PM.

Operations put the TeV into shot setup at 9:14 PM.

At 10:07 PM, the TeV suffered a sector F3 quench when beam was injected onto other beam due to a sequencer change. The TeV had recovered from the quench by the end of the shift.

Friday June 10
The midnight shift began with Operations monitoring the stash, with Pbar stacking and working on their beam cooling, with the TeV in a dry squeeze, and with MiniBooNE and NuMI taking beam.

Experts had the NuMI kicker repaired and tested by 12:53 AM.

At 1:47 AM, a NuMI VSA problem arose that required an expert. He came in and around 4 AM reported that a beam-cooling problem was affecting the RF.

Operations put the TeV through a dry squeeze at 3:32 AM.

At 4:57 AM, the MI BPMs stopped reading data. Operators called in several experts.

The Plan for Friday
The plans for today and this weekend are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/10/2005   email Fermilab